Triangle ABC is an isoceles triangle. Its base is 16.2 cm and one base angle is 63 degrees 20 minutes. Find the length of one of the congruent sides to the nearest hundredth of a cm.
Thanks

Answer by lwsshak3(11628)

Triangle ABC is an isoceles triangle. Its base is 16.2 cm and one base angle is 63 degrees 20 minutes. Find the length of one of the congruent sides to the nearest hundredth of a cm.
Thanks
..
63 deg 20 min=63.3 deg (rounded)
The angle opposite the base=180-63.3-63.3=53.4
Let x=one of the congruent sides
Using Law of sines,
Sin53.4/16.2=sin 63.3/x
x=(sin 63.3/sin 53.4)*16.2
x=18.03 cm
ans:
length of congruent side=18.03 cm